OpenStudy (anonymous):

What are the factors of 3x2 + 10x + 3

OpenStudy (anonymous):

use the quadratic formula to solve this one

pooja195 (pooja195):

@Serenaishurting you can also use factoring is you wish

pooja195 (pooja195):

*if

OpenStudy (anonymous):

how?

pooja195 (pooja195):

Does it specify which method to use?

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ }\] where \[ax^2+bx+c=0\]

OpenStudy (anonymous):

no it doesnt

pooja195 (pooja195):

ok so you can use the formula then since @saseal typed it out ;p \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ }\] where \[ax^2+bx+c=0\]

pooja195 (pooja195):

Start by listing your abc values

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\]

OpenStudy (anonymous):

sorry lagged and the 2a didnt show up

OpenStudy (anonymous):

okay thanks i got it.

OpenStudy (anonymous):

yw

OpenStudy (radar):

Another way is possible.....just factor it by noting that both the coefficient of the squared term "3" and the numeric constant is "3" both are prime numbers. This means less trial and error.

OpenStudy (radar):

It factors quite nicely: (x + 3) (3x + 1)

OpenStudy (anonymous):

can yall help with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Graph of a cubic polynomial that falls to the left and rises to the right with x intercepts negative 2, negative 1, and 2 Which of the following functions best represents the graph?

OpenStudy (anonymous):

Which of the following functions best represents the graph? f(x) = x3 + x2 − 4x − 4 f(x) = x3 + 4x2 − x − 4 f(x) = x3 + 3x2 − 4x − 12 f(x) = x3 + 2x2 − 4x − 8

OpenStudy (anonymous):

your best bet is to graph it out, cubic stuff is pretty messy

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lemme try teaching you how to factor this without graphing

OpenStudy (anonymous):

okay

OpenStudy (radar):

This problem is providing you the factors. Just multiply them and get the answer.

OpenStudy (anonymous):

we take this 1st option\[f(x)=x^3+x^2-4x-4\] we can rewrite this as \[f(x)=x^2(x+1)-4(x-1)\]

OpenStudy (radar):